For the next fiscal year, you forecast net income of $50,000 and ending assests of $500,000. Your firm's payout ratio is 10%. Your beginning stockholders equity is $300,000 and your beginning total liabilities are $120,000. Your non-debt liabilites such as accounts payable are forecasted to increase by $10,000. What is you net new financing needed for next year?
1
Expert's answer
2013-02-28T11:46:01-0500
Additions to equity = net income x retention ratio= 50,000 x (1-.10) = 45,000
Dividents per share = 300,000*10% = 30,000
Total capital = 45,000 + 30,000 + 50,000 + 300,000 = 425,000
Capital + Total Liabilities = 425,000 + 120,000 + 10,000 = 555,000
So, you need additional 55,000 of net new financing.